OpenStudy (anonymous):

A salesperson earns $500 per month plus a commission of 20% of sales. Describe the sales needed to receive a total income that exceeds $3200 a month

OpenStudy (anonymous):

500 + 0.2x = 3200 0.2x = 2700 x = 2700 / 0.2 x = 13500 We can check by plugging in 13500 So we have 500 + 0.2(13500) = 3200 500 + 2700 = 3200 3200 = 3200 So the answer is any amount of over 13500 will result in more than $3200 earned.

OpenStudy (anonymous):

THANK YOU!!